# Graph Attention Network-based Block Propagation with Optimal AoI and Reputation in Web 3.0

### v1

@@ Abstract
Web 3.0 is recognized as a pioneering paradigm that empowers users to securely oversee data without reliance on a centralized authority. Blockchains, as a core technology to realize Web 3.0, can facilitate decentralized and transparent data management. Nevertheless, the evolution of blockchain-enabled Web 3.0 is still in its nascent phase, grappling with challenges such as ensuring efficiency and reliability to enhance block propagation performance. In this paper, we design a Graph Attention Network (GAT)-based reliable block propagation optimization framework for blockchain-enabled Web 3.0. We first innovatively apply a data-freshness metric called age of information to measure block propagation efficiency in public blockchains. To achieve the reliability of block propagation, we introduce a reputation mechanism based on the subjective logic model, including the local and recommended opinions to calculate the miner reputation value. Moreover, considering that the GAT possesses the excellent ability to process graph-structured data, we utilize the GAT with reinforcement learning to obtain the optimal block propagation trajectory. Numerical results demonstrate that the proposed scheme exhibits the most outstanding block propagation efficiency and reliability compared with traditional routing algorithms.
@@
Web 3.0, block propagation, age of information, reputation, graph attention network.
@@ I Introduction
With the incredible ability to empower secure, transparent, and tamper-proof data transactions through decentralized ledgers, blockchains have attracted significant attention from both academia and industry[3]. Based on consensus algorithms and encryption technologies, blockchains possess the ability to facilitate the creation of cryptocurrencies and drive the development of decentralized applications and smart contracts, assuming a pivotal role across various domains, such as decentralized finance[4], Metaverse[5], and Internet of Vehicles[6]. In Web 3.0, a considerable volume of transactions involving digital products, such as Non-Fungible Tokens, transpires among users. This necessitates the adoption of blockchain technologies to ensure secure storage and efficient management of these transactions[2]. Therefore, blockchains are considered fundamental and indispensable technologies for the advancement of Web 3.0.
@@
To address the above challenges, in this paper, we design a Graph Attention Network (GAT)-based reliable block propagation optimization framework for blockchain-enabled Web 3.0. Specifically, since Age of Information (AoI) as a well-established metric can capture information freshness[9], we apply the AoI to measure block propagation efficiency in public blockchains. Then, to achieve reliable block propagation optimization, we propose a miner reputation mechanism based on the subjective logic model[10], including the local and recommended opinions to measure miner reliability in public blockchains. Furthermore, we utilize the GAT with Reinforcement Learning (RL) to extract the relational presentation between each miner and its adjacent miner in a graph-structured miner network, thereby facilitating the formulation of the miner selection process. By integrating the aforementioned methods, the framework can ultimately derive the optimal block propagation trajectory. The contributions of this paper are summarized as follows:
  • We adopt the AoI as a data-freshness metric to measure block propagation efficiency in public blockchains, which comprehensively considers the procedures of block waiting and block propagation.
  • To ensure the reliability of block propagation, we propose a miner reputation mechanism based on the subjective logic model, where the subjective logic model as a widely adopted mathematical tool utilizes the local and recommended miner opinions to effectively calculate miner reputation values in the miner network.
  • To achieve reliable block propagation optimization, we propose a GAT-based block propagation optimization model, in which we design a novel GAT-based architecture with RL to minimize the overall AoI of block propagation given the reputation constraint, thereby obtaining the optimal block propagation trajectory.
  • Compared with Greedy and Gossip algorithms, our model exhibits outstanding performance and achieves the most efficient and reliable block propagation. For example, when the number of miners is set to $49$, the overall AoI of our model is $2.4\%-22.6\%$ lower, while the total value of miner reputations of our model is $7.4\%-34.77\%$ higher than that of Greedy and Gossip algorithms.
@@
The remainder of the paper is organized as follows: Section II reviews related literature. In Section III, we propose the system model including the block propagation mechanism and the proposed framework. In Section IV, we formulate the problem of minimizing the overall AoI of block propagation with the given reputation constraint. In Section V, we introduce the GAT with the encoder-decoder architecture. In Section VI, we introduce RL with a rollout baseline to train the network. Section VII presents the numerical results of the proposed schemes. Finally, the paper concludes with Section VIII.
@@ II Related Work
$\Delta$ Average AoI over the time range (0, $\delta$)
@@
Web 3.0, also regarded as the semantic web, represents the cutting edge of web development, driven by Artificial Intelligence (AI), Internet of Things, and blockchain technologies. Nowadays, advancing the security of data storage and enhancing the efficiency of information interaction have remained important themes in Web research. Therefore, blockchain-enabled Web 3.0 has attracted significant attention, where Web 3.0 leverages blockchain technologies for the transparent, traceable, and decentralized recording of on-chain content. Some efforts have been conducted to achieve blockchain-enabled Web 3.0[2, 8, 11]. For example, the authors in[2] introduced a Web 3.0 framework powered by quantum blockchain technologies and discussed potential challenges and applications of integrating quantum blockchain in Web 3.0[2]. The authors in [8] proposed a semantic exchange framework leveraging blockchain technologies and discussed the relationship between semantic communication and blockchain for Web 3.0[8]. The authors in[11] developed an innovative architecture for Web 3.0, integrating AI and blockchain to establish a metaverse-based framework that aims to tackle the existing privacy and security challenges of the current Web 2.0 while improving user experience and data control. Moreover, the authors emphasized the importance of blockchain technologies in the construction of Web 3.0[11]. However, the majority of existing works do not delve into the optimization of blockchain performance as a means to enhance Web 3.0 performance. Therefore, it is important to optimize blockchain performance for Web 3.0, especially optimizing block propagation.
@@ II-C Block Propagation Optimization
Motivated by the aforementioned research gaps, we propose a GAT-based reliable block propagation optimization framework for the performance enhancement of blockchain-enabled Web 3.0, in which we apply the AoI to quantify block propagation efficiency and propose a miner reputation mechanism to ensure the reliability of block propagation.
@@ III-A Block Propagation Mechanism
According to [21], we introduce a block propagation mechanism where the new block is propagated to a single miner, departing from the flooding propagation characteristic of the traditional gossip protocol. The detailed process is depicted in part A of Fig. 1. To prevent redundant transmissions to miners that have already received the new block, the new block is not forwarded directly[21]. Instead, a miner would send an inv message containing the hash value of the new block to its adjacent miner before propagating the block. If the adjacent miner has never received the block, the adjacent miner will issue a getdata message to request the new block [21, 22], and the getdata message is then queued until the block has been received and the difficulty check has been done[21]. To speed up block propagation, the adjacent miner sending the getdata message also sends an inv message to its adjacent miner. Additionally, the block verification process is divided into two phases, i.e., the block difficulty check and transaction validation [21]. Once the block difficulty check is completed, the block is sent out without further transaction validation, aiming to minimize processing time [21].
@@ III-B Framework Design
As shown in Fig. 1, the proposed GAT-based reliable block propagation optimization framework for blockchain-enabled Web 3.0 is structured into three layers, i.e., a physical layer, a data layer, and a network layer. Specifically, the first layer, functioning as the foundation of the proposed framework, is the physical layer, which exists in various applications in Web 3.0, such as decentralized autonomous organizations and decentralized marketplaces for digital products[1]. The second layer is the data layer, where blockchains provide secure and reliable data management for Web 3.0 applications[2]. The final layer is the network layer, which is exacted from the blockchains of the data layer. Especially, the network layer introduces the proposed framework design for achieving block propagation optimization, including four parts. Part A presents a block propagation protocol that shows how a new block propagates between miners in the miner network[21]. Part B introduces the AoI utilized to measure the block propagation efficiency in public blockchains. Part C focuses on the miner reputation mechanism, which utilizes the subjective logic model to calculate miner reputations for ensuring reliable block propagation[10]. Part D shows the GAT, which works directly on graph-structured data and possesses the excellent ability to capture the relationships between miners in the miner network[23]. The proposed framework aims to achieve efficient and reliable block propagation optimization, thereby enhancing the overall performance of Web 3.0 systems.
@@ IV Problem Formulation for Efficient and Reliable Block Propagation
In this section, we first introduce the AoI to quantify block propagation efficiency in public blockchains. Then, we propose a miner reputation mechanism based on the subjective logic model. In this paper, we consider that there is a set =$\mathcal{M}=\left\{1,\ldots,i,\ldots,j,\ldots,M\right\}$ of $M$ miners with the nature of random mobility in the miner network, and the block propagation process can be considered an M/M/1 system following a first-come-first-served policy.
IV-A Age of Information for Block Propagation
AoI as an effective data-freshness metric, has been widely utilized in remote system status, which is denoted as the elapsed time from the generation of the latest received status update[24]. Motivated by [24, 25], in this paper, we utilize the AoI to quantify block propagation efficiency, which is denoted as the time elapsed from a getdata message sent by a miner to its successful receipt of the new block, consisting of procedures of block waiting and block propagation[25]. Especially, low AoI indicates enhanced block propagation efficiency, which reduces the probability of block forks and insufficient signature collection[26], consequently avoiding the occurrence of serious system malfunction of Web 3.0. We denote $t_{i}$ as the time that miner $i$ sends a getdata message and $t_{i}^{{}^{\prime}}$ as the time that miner $i$ successfully receives the block. Thus, the instantaneous AoI of block propagation between adjacent miners at any time $t$ is expressed as [24]
@@ IV-A Age of Information for Block Propagation
where $I(t)=\max\{i|t_{i}^{{}^{\prime}}\leq t,i\in\mathcal{M}\}$ represents the index of the miner that recently received the block. Based on (1), the overall AoI over the time range $(0,\delta)$ is given by
@@
and the average AoI over the time range $(0,{\delta})$ is expressed as
@@
Now, we show the characterization of the average AoI of block propagation between adjacent miners in public blockchains. We define the $i$-th time interval, denoted by $X_{i}$, as the time elapsed from miner $(i-1)$ sending the getdata message to miner $i$ sending the getdata message, which is given by
@@
Therefore, the AoI of block propagation between miner $(i-1)$ and miner $i$ in public blockchains is given by[24]
@@
$\Gamma_{i}=\frac{B_{block}}{W\log_{2}\Big{(}1+\frac{\rho_{s}c^{0}d_{i-1,i}^{-\varepsilon}}{N_{0}W}\Big{)}},$ (11)
where $B_{block}$, $W$, $\rho_{s}$, ${d_{i-1,i}}$, $c^{0}$, $\varepsilon$, and $N_{0}$ represent the block size, the link bandwidth between adjacent miners, the transmit power of miners, the distance between miner $(i-1)$ and miner $i$, the unit channel power gain, the path-loss coefficient, and the noise power density, respectively. For the system time $D_{i}$ of miner $i$, $f_{D_{i}}(t)$ is given by[24]
@@
Therefore, based on (7), the AoI between miner $(i-1)$ and miner $i$ can be given by
@@ IV-B Reliable Block Propagation based on the Subjective Logic Model
The subjective logic, utilized for the fusion of probabilistic information[10], is an effective tool to quantify the reliability of miners in the blockchain network[28], which utilizes logic and recommended opinions to express the subjective beliefs[29]. Based on the subjective logic model, both direct and indirect reputations from other adjacent miners are considered to derive reputation opinions of miners[28], and the subjective logic divides miner reputations into three levels, i.e., trust, distrust, and uncertainty[28, 10]. For the miner $j$, its compositive reputation opinions consist of the indirect reputation opinions of other adjacent miners and the direct reputation opinions of miner $i$ that directly interact with the miner $j$. Especially, the direct reputation opinions of miner $i$ are considered to be local reputation opinions, and the indirect reputation opinions of other adjacent miners are considered to be recommended reputation opinions[28].
@@ IV-B2 Recommended opinions for subjective logic
For recommender $s$ and miner $j$, the interaction frequency is defined as the ratio of the interactions between recommender $s$ and miner $j$ to the average value of interactions between recommender $s$ and miners, which is denoted as $IF_{s\to j}=H_{s\to j}/\overline{H}_{s}$, where $H_{s\to j}=\alpha_{s\to j}+\beta_{s\to j}$ and $\overline{H}_{s}=\left(\frac{1}{M}\right)\sum_{j=1}^{M}H_{s\to j}$[30]. Furthermore, the weight of recommended opinions of recommender $s$ is defined as $\omega_{s\to j}=\delta_{s\to j}\cdot IF_{s\to j}$, where $\delta_{s\to j}$ is a pre-defined coefficient for the calculation of the recommended reputation. Therefore, the recommended reputation opinion of recommender $s$ to miner $j$ can be obtained as

### v2

# Graph Attention Network-based Block Propagation with Optimal AoB and Reputation in Web 3.0

@@ Abstract
Web 3.0 is recognized as a pioneering paradigm that empowers users to securely oversee data without reliance on a centralized authority. Blockchains, as a core technology to realize Web 3.0, can facilitate decentralized and transparent data management. Nevertheless, the evolution of blockchain-enabled Web 3.0 is still in its nascent phase, grappling with challenges such as ensuring efficiency and reliability to enhance block propagation performance. In this paper, we design a Graph Attention Network (GAT)-based reliable block propagation optimization framework for blockchain-enabled Web 3.0. We first innovatively apply a data-freshness metric called age of block to measure block propagation efficiency in public blockchains. To achieve the reliability of block propagation, we introduce a reputation mechanism based on the subjective logic model, including the local and recommended opinions to calculate the miner reputation value. Moreover, considering that the GAT possesses the excellent ability to process graph-structured data, we utilize the GAT with reinforcement learning to obtain the optimal block propagation trajectory. Numerical results demonstrate that the proposed scheme exhibits the most outstanding block propagation efficiency and reliability compared with traditional routing mechanisms.
@@
Web 3.0, block propagation, age of block, reputation, graph attention network.
@@ I Introduction
With the incredible ability to empower secure, transparent, and tamper-proof data transactions through decentralized ledgers, blockchains have attracted significant attention from both academia and industry[3]. Based on consensus mechanisms and encryption technologies, blockchains possess the ability to facilitate the creation of cryptocurrencies and drive the development of decentralized applications and smart contracts, assuming a pivotal role across various domains, such as decentralized finance[4], Metaverse[5], and Internet of Vehicles[6]. In Web 3.0, a significant volume of transactions involving digital products, such as Non-Fungible Tokens, transpires among users, which poses challenges to data security. To this end, blockchain-enabled Web 3.0 emerges as a necessary strategy to ensure secure storage and efficient management of these transactions[2]. Furthermore, the blockchain-enabled Web 3.0 facilitates decentralized data exchange and sharing, which enables participants to circumvent the dominance of tech giants and gain control over user-generated content[7]. Therefore, blockchains are widely regarded as fundamental and indispensable technologies for the progress of Web 3.0.
@@
To address the above challenges, in this paper, we design a Graph Attention Network (GAT)-based reliable block propagation optimization framework for blockchain-enabled Web 3.0. Specifically, since Age of Information (AoI) as a well-established metric can capture information freshness[9], we apply the Age of Block (AoB), which is similar to AoI, to measure block propagation efficiency in public blockchains. Then, to achieve reliable block propagation optimization, we propose a miner reputation mechanism based on the subjective logic model[10], including the local and recommended opinions to measure miner reliability in public blockchains. Furthermore, we utilize the GAT with Reinforcement Learning (RL) to extract the relational presentation between each miner and its adjacent miner in a graph-structured miner network, thereby facilitating the formulation of the miner selection process. By integrating the aforementioned methods, the framework can ultimately derive the optimal block propagation trajectory. The contributions of this paper are summarized as follows:
  • We adopt the AoB as a data-freshness metric to measure block propagation efficiency in public blockchains, which has a specific consideration of the block consensus mechanism. For example, the proposed AoB innovatively provides a comprehensive examination of the procedures of getdata message arrival time interval, block waiting time, and block propagation time. Moreover, unlike [11], we infer the specific formulas considering the characteristics of block propagation in public blockchains.
  • To ensure the reliability of block propagation, we propose a miner reputation mechanism based on the subjective logic model, where the subjective logic model as a widely adopted mathematical tool utilizes the local and recommended miner opinions to effectively calculate miner reputation values in the miner network. Moreover, the miner reputation mechanism has been specifically designed in both local and recommended opinions to improve the reputation calculation.
  • To achieve reliable block propagation optimization, we propose a GAT-based block propagation optimization model to delve into more intricate aspects during the block consensus process than existing consensus mechanisms, in which we design a novel GAT-based architecture with RL to minimize the overall AoB of block propagation given the reputation constraint, thereby obtaining the optimal block propagation trajectory. Moreover, considering the consensus mechanism of public blockchains (i.e., Proof of Work), the optimization conducted by the GAT model is based on a specific number of miners.
  • Compared with Greedy and Gossip mechanisms, our model exhibits outstanding performance and achieves the most efficient and reliable block propagation. For example, when the number of miners is set to $49$, the overall AoB of our model is $2.4\%-22.6\%$ lower than that of Greedy and Gossip mechanisms. Moreover, the total value of miner reputations of our model is $7.4\%-34.77\%$ higher than that of Greedy and Gossip mechanisms.
@@
The remainder of the paper is organized as follows: Section II reviews related literature. In Section III, we propose the system model including the block propagation mechanism and the proposed framework. In Section IV, we formulate the problem of minimizing the overall AoB of block propagation with the given reputation constraint. In Section V, we introduce the GAT with the encoder-decoder architecture. In Section VI, we introduce RL with a rollout baseline to train the network. Section VII presents the numerical results of the proposed schemes. Finally, the paper concludes with Section VIII.
@@ II Related Work
$\Delta$ Average AoB over the time range (0, $\delta$)
@@
Web 3.0, also regarded as the semantic web, represents the cutting edge of web development, driven by Artificial Intelligence (AI), Internet of Things, and blockchain technologies. Nowadays, the intertwining of blockchain technologies with Web 3.0 forms a symbiotic relationship, giving rise to blockchain-enabled Web 3.0, which is effective for the transparent, traceable, and decentralized recording of on-chain content. Therefore, research aimed at enhancing blockchain performance in Web 3.0 holds unique significance, serving as a key motivation for our work. Some efforts have been conducted to achieve blockchain-enabled Web 3.0[2, 7, 12]. For example, the authors in[2] introduced a Web 3.0 framework powered by quantum blockchain technologies and discussed potential challenges and applications of integrating quantum blockchain in Web 3.0[2]. The authors in [7] proposed a semantic exchange framework leveraging blockchain technologies and discussed the relationship between semantic communication and blockchain for Web 3.0[7]. The authors in[12] developed an innovative architecture for Web 3.0, integrating AI and blockchain to establish a metaverse-based framework that aims to tackle the existing privacy and security challenges of the current Web 2.0 while improving user experience and data control. Moreover, the authors emphasized the importance of blockchain technologies in the construction of Web 3.0[12]. However, the majority of existing works do not delve into the optimization of blockchain performance as a means to enhance Web 3.0 performance. Therefore, it is important to optimize blockchain performance for Web 3.0, especially optimizing block propagation.
@@ II-C Block Propagation Optimization
Motivated by the aforementioned research gaps, we propose a GAT-based reliable block propagation optimization framework for the performance enhancement of blockchain-enabled Web 3.0, in which we apply the AoB to quantify block propagation efficiency and propose a miner reputation mechanism to ensure the reliability of block propagation.
@@ III-A Block Propagation Mechanism
For the block propagation in public blockchains, we introduce the block propagation mechanism which is similar to the traditional gossip protocol[22, 23]. However, for the optimization of specific block propagation trajectories, the introduced block propagation mechanism directs the propagation of a new block to a single miner, departing from the flooding propagation characteristic of the gossip. The detailed process is depicted in Part A of Fig. 1. To prevent redundant transmissions to miners that have already received the new block, the new block is not forwarded directly[22]. Instead, a miner would send an inv message containing the hash value of the new block to its adjacent miner before propagating the block. If the adjacent miner has never received the block, the adjacent miner will issue a getdata message to request the new block [22, 24], and the getdata message is then queued until the block has been received and the difficulty check has been done[22]. To speed up block propagation, the adjacent miner sending the getdata message also sends an inv message to its adjacent miner. Additionally, the block verification process is divided into two phases, i.e., the block difficulty check and transaction validation [22]. Once the block difficulty check is completed, the block is sent out without further transaction validation, aiming to minimize processing time [22].
@@ III-B Framework Design
As shown in Fig. 1, the proposed GAT-based reliable block propagation optimization framework for blockchain-enabled Web 3.0 is structured into three layers, i.e., a physical layer, a data layer, and a network layer. Specifically, the first layer, functioning as the foundation of the proposed framework, is the physical layer, which exists in various applications in Web 3.0, such as decentralized autonomous organizations and decentralized marketplaces for digital products[1]. The second layer is the data layer, where blockchains provide secure and reliable data management for Web 3.0 applications[2]. The final layer is the network layer, which is exacted from the blockchains of the data layer. Especially, the network layer introduces the proposed framework design for achieving block propagation optimization, including four parts. Part A presents a block propagation protocol that shows how a new block propagates between miners in the miner network[22]. Part B introduces the AoB utilized to measure the block propagation efficiency in public blockchains. Part C focuses on the miner reputation mechanism, which utilizes the subjective logic model to calculate miner reputations for ensuring reliable block propagation[10]. Part D shows the GAT, which works directly on graph-structured data and possesses the excellent ability to capture the relationships between miners in the miner network[25]. The proposed framework aims to achieve efficient and reliable block propagation optimization, thereby enhancing the overall performance of Web 3.0 systems.
@@ IV Problem Formulation for Efficient and Reliable Block Propagation
In this section, we first introduce the AoB to quantify block propagation efficiency in public blockchains. Then, we propose a miner reputation mechanism based on the subjective logic model. In this paper, we consider that there is a set =$\mathcal{M}=\left\{1,\ldots,i,\ldots,j,\ldots,M\right\}$ of $M$ miners with the nature of random mobility in the miner network, and the block propagation process can be considered an M/M/1 system following a first-come-first-served policy.
IV-A Age of Block for Block Propagation
Since AoI as an effective data-freshness metric, has been widely utilized in remote system status, which is denoted as the elapsed time from the generation of the latest received status update[26]. In this paper, we utilize the AoB to quantify block propagation efficiency motivated by [26, 11], which is denoted as the time elapsed from a getdata message sent by a miner to its successful receipt of the new block, consisting of procedures of getdata message arrival time interval, block waiting time, and block propagation time[11]. We denote $t_{i}$ as the time that miner $i$ sends a getdata message and $t_{i}^{{}^{\prime}}$ as the time that miner $i$ successfully receives the block. Thus, the instantaneous AoB of block propagation between adjacent miners at any time $t$ is expressed as [26]
@@ IV-A Age of Block for Block Propagation
where $I(t)=\max\{i|t_{i}^{{}^{\prime}}\leq t,i\in\mathcal{M}\}$ represents the index of the miner that recently received the block. Based on (1), the overall AoB over the time range $(0,\delta)$ is given by
@@
and the average AoB over the time range $(0,{\delta})$ is expressed as
@@
Now, we show the characterization of the average AoB of block propagation between adjacent miners in public blockchains. We define the $i$-th time interval, denoted by $X_{i}$, as the time elapsed from miner $(i-1)$ sending the getdata message to miner $i$ sending the getdata message, which is given by
@@
Therefore, the AoB of block propagation between miner $(i-1)$ and miner $i$ in public blockchains is given by[26]
@@
$\Gamma_{i}=\frac{B_{block}}{b\log_{2}\Big{(}1+\frac{\rho_{s}c^{0}d_{i-1,i}^{-\varepsilon}}{N_{0}b}\Big{)}},$ (12)
where $B_{block}$, $b$, $\rho_{s}$, ${d_{i-1,i}}$, $c^{0}$, $\varepsilon$, and $N_{0}$ represent the block size, the channel bandwidth between adjacent miners, the transmit power of miners, the distance between miner $(i-1)$ and miner $i$, the unit channel power gain, the path-loss coefficient, and the noise power density, respectively. For the system time $D_{i}$ of miner $i$, $f_{D_{i}}(t)$ is given by[26]
@@
Therefore, based on (7), the AoB between miner $(i-1)$ and miner $i$ can be given by
@@
(15) demonstrates that low overall AoB means the comprehensive reduction of the time of block propagation process, including the getdata message arrival time interval, the block waiting time, and the block propagation time, which indicates the enhanced block propagation efficiency. Furthermore, the PDF of a block to be mined can be given by[28]
$f(t;\mu)=\mu e^{-\mu t}.$ (16)
Based on (16), we can obtain the block fork probability, which is a function of block propagation time and the average getdata delivery ratio of miners, given by [28]
$F(t)=P(X_{i}\leq{\Gamma})=\int_{0}^{{\Gamma}}f(t)\mathrm{d}t=1-e^{-\mu{{\Gamma}}},$ (17)
where $\Gamma$ is the overall block propagation time. Thus, combining (17) and (26), it can be inferred that a low overall AoB can reduce the block fork probability when $\mu$ is a fixed value.
@@ IV-B Reliable Block Propagation based on the Subjective Logic Model
The subjective logic model, which utilizes logic and recommended opinions to express subjective beliefs, is a comprehensive model of subjective beliefs evaluation. Moreover, the miner interactions in the blockchain can be considered as a form of group behavior, which can be calculated by direct and indirect opinions[10]. Therefore, for a better fusion of probabilistic information, we utilize the subjective logic model as an effective tool to quantify the reliability of miners in the blockchain network.
Based on the subjective logic model, the subjective logic divides miner reputations into three levels, i.e., trust, distrust, and uncertainty[29, 10]. For the miner $j$, its compositive reputation opinions consist of the indirect reputation opinions of other adjacent miners and the direct reputation opinions of miner $i$ that directly interact with the miner $j$. Especially, the direct reputation opinions of miner $i$ are considered to be local reputation opinions, and the indirect reputation opinions of other adjacent miners are considered to be recommended reputation opinions[29].
@@ IV-B2 Recommended opinions for subjective logic
Especially, considering that miners who engage in more positive interactions should obtain a larger proportion of recommendations, we introduce $\gamma^{rec}\in(1,\infty)$ as an enhancement factor reflecting the influence weight of positive interactions on recommendations. Moreover, for recommender $s$ and miner $j$, the interaction frequency is defined as the ratio of the interactions between recommender $s$ and miner $j$ to the average value of interactions between recommender $s$ and miners, which is denoted as $IF_{s\to j}=H_{s\to j}/\overline{H}_{s}$, where $H_{s\to j}=\gamma^{rec}\alpha_{s\to j}+\beta_{s\to j}$ and $\overline{H}_{s}=\left(\frac{1}{M}\right)\sum_{j=1}^{M}H_{s\to j}$ [30]. Furthermore, the weight of recommended opinions of recommender $s$ is defined as $\omega_{s\to j}=\delta_{s\to j}\cdot IF_{s\to j}$, where $\delta_{s\to j}$ is a pre-defined coefficient for the calculation of the recommended reputation. Therefore, the recommended reputation opinion of recommender $s$ to miner $j$ can be obtained as